OpenStudy (anonymous):

Eliminate the parameter to write the parametric equations as a rectangles equation. X=(1)/(t-2). Y=4t+5

OpenStudy (anonymous):

I'm suppose to get t by itself in the x equation and just plug it into the y equation

OpenStudy (faiqraees):

Write the first equation making t the subject

OpenStudy (anonymous):

Yeah that's what I don't understand

OpenStudy (faiqraees):

Multiply both sides by t-2 and then divide both sides by X of the first equation

OpenStudy (anonymous):

Now I'm left with t-2=1/x

OpenStudy (faiqraees):

now add 2 on both sides

OpenStudy (anonymous):

T=(1/x)+2

OpenStudy (anonymous):

?

OpenStudy (faiqraees):

Now substitute the value of t in the second equation

OpenStudy (anonymous):

Y=4(1/x + 2)+5

OpenStudy (faiqraees):

now simplify

OpenStudy (anonymous):

Y=4/x + 13

OpenStudy (faiqraees):

you've the answer btw more appropriate form would be after multiplying the equation by x

OpenStudy (anonymous):

What would the answer look then

OpenStudy (faiqraees):

yx=4x+13

OpenStudy (anonymous):

Can we do a few more. I just need to make sure they are right as I have already finished them

OpenStudy (faiqraees):

sure but one question on each post

OpenStudy (anonymous):

So I need a new post?

OpenStudy (faiqraees):

yes

OpenStudy (anonymous):

Will do